31x – 8(3x – 0.745) = 6x + 0.205 Solve the equation.
Answer by ptaylor(1894)

First, multiply to get rid of the parens:
31x-24x+5.96=6x+0.205 subtract 6x and also 5.96 from both sides
31x-24x+5.96-5.96-6x=0.205-5.96+6x-6x collecting like terms
x=-5.755
ck
substitute x=-5.755 into the original equation:
31(-5.755)-8(3(-5.755)-0.745)=6(-5.755)+0.205 and we get
-178.405-8(-17.265-0.745)=-34.53+0.205 simplifying
-178.405+144.08=-34.325
-34.325=-34.325


Hope this helps----ptaylor